# First investigation of void statistics in numerical relativity simulations

### v1

@@ 4.2 Void fraction
Voids comprise the majority of the volume of the late-time Universe. Constraining the precise fraction of voids is important for the Timescape model of cosmic backreaction (Wiltshire, 2007a, b, 2009; Duley et al., 2013), since it is the free parameter which replaces $\Omega_{\rm M0}$ in standard FLRW models.
@@ 4.3 Finding void centres
We use $f(\rho)=\rho_{\mathrm{ave}}-\rho$ to ensure that the centre is weighted towards areas of low density. When matter is distributed unevenly around a void, its centre of mass will differ from its centre of volume. Under this definition, the decision to include or exclude a higher-density cell has less impact on the centre that it would have on the centre of volume. Thus, it may be more robust to the choice of exclusion threshold or slight changes in the density near it, which may cause the set of excluded cells to change.
In the 4 $h^{-1}$ Mpc resolution simulation, the mean distance between a void’s centre of mass $\mathbf{X}^{\mathrm{cm}}_{v}$ and its centre of volume $\mathbf{X}^{\mathrm{bc}}_{v}$ is approximately 0.88 $h^{-1}$ Mpc, or 0.22 grid cells, and has an upper bound of a little over one cell. Since this difference is small with respect to our grid resolution, we consider both to be an appropriates measures of centre in our case. We use the centre of volume $\mathbf{X}^{\mathrm{bc}}_{v}$ hereafter for the sake of comparison with existing literature.
@@ 4.4.1 Constructing stacked radial profiles
Figure 3: Upper: The stacked density contrast (solid curve) on a spherical shell at distance $r$ from the stacked void centre in the 4 $h^{-1}$ Mpc simulation, with each void individually normalised by its effective radius $R_{\mathrm{eff}}$. The shaded areas show the middle  68.5% and  95% of $\delta$ values for cells on that shell. Lower: The fraction of cells marked as void cells as a function of radius (solid) compared to perfect spheres (dotted) and a set of gridded spheres with radius 3–5 grid cells (dashed).
Figure 4: Upper: The stacked density contrast (solid curve) on a spherical shell at distance $r$ from the stacked void centre in the 12 $h^{-1}$ Mpc simulation, with each void individually normalised by its effective radius $R_{\mathrm{eff}}$. The shaded areas show the middle  68.5% and  95% of $\delta$ values for cells on that shell. Lower: The fraction of cells marked as void cells as a function of radius (solid) compared to perfect spheres (dotted) and a set of gridded spheres with radius 3–5 grid cells (dashed).
@@
Because the steps are in terms of a fraction of each void’s radius, individual profiles can easily be stacked and averaged to produce a radial profile for a sample of voids. The process of stacking a large number of voids produces a spherically symmetric object, despite individual voids not being necessarily spherical. Stacking voids in this manner has other uses beyond extracting radial profiles. It can also be used to amplify the signal of the integrated Sachs–Wolfe effect (Granett et al., 2008), as temperature fluctuations in the CMB are larger than the temperature change due to these effects in any single structure; and for the Alcock–Paczyński (1979) test (Sutter et al., 2014b; Hamaus et al., 2020).
@@ 4.4.2 Stacked radial density profiles
We stack 2000 randomly selected voids (only a subset of the more than 30000 voids per sample, due to computational limitations) to produce averaged radial profiles for the density contrast in the voids. The top panels of Fig. 3 and Fig. 4 show these profiles, normalised by each void’s effective radius, for the 4 $h^{-1}$ Mpc and 12 $h^{-1}$ Mpc resolution simulations, respectively. The shaded areas of these figures show the density contrast of the middle  68.5% and  95% of cells at each rescaled radial distance. These shaded bands show that walls begin appearing in some directions even when $r<R_{\mathrm{eff}}$, where the average density contrast is still negative. Likewise, the overcompensation continues much further away from the centre than $R_{\mathrm{eff}}$, continuing to be visible past $r=2R_{\mathrm{eff}}$. This is a natural consequence of the fact that the voids are not all close to spherical. The lower panels in these figures show this explicitly, showing what proportion of points at a given distance are inside of the void or outside of it (solid curve). For comparison, these panels also show the case of spheres at infinite resolution (dashed line; a step function) and a set of spheres of radius 3–5 grid cells (dashed curve) – comparable to the effective radius of typical voids in these simulations. The latter is not a step function, due to the spheres being represented on the finite-resolution grid.
@@
Figure 5: The stacked radial density contrast from Fig. 3 (solid blue; 4 $h^{-1}$ Mpc simulation) and its fitted HSW profile (dashed black), and the stacked radial density contrast from Fig. 4 (dash-dotted orange; 12 $h^{-1}$ Mpc simulation) and its fitted HSW profile (dotted black). These HSW profiles use (8), fitting only the typical four parameters.
Figure 6: Upper: The stacked radial density contrast from Fig. 3 (solid blue; 4 $h^{-1}$ Mpc simulation) and its fitted HSW profile (dashed black), and the stacked radial density contrast from Fig. 4 (dash-dotted orange; 12 $h^{-1}$ Mpc simulation) and its fitted HSW profile (dotted black). These HSW profiles use (9), with the fifth fitted parameter $R_{\mathrm{v}}$ in place of $R_{\mathrm{eff}}$ to determine the profile. The $x$-axis is normalised by the original $R_{\mathrm{eff}}$. Lower: Difference between HSW density contrast and stacked radial density contrast for the 4 $h^{-1}$ Mpc resolution simulation (solid) and the 12 $h^{-1}$ Mpc resolution simulation (dash-dotted).
@@ 4.4.3 Fitting HSW profiles
where $R_{\mathrm{v}}$ is a parameter that determines the size of the void in place of $R_{\mathrm{eff}}$. The rows of Table 1 with “fitting $R_{\mathrm{v}}$” list the adjusted parameter sets. The fitted $R_{\mathrm{v}}$ is 25% and 32% larger than the effective radius for 4 $h^{-1}$ Mpc and 12 $h^{-1}$ Mpc resolution, respectively. Fig. 6 shows the HSW profiles with these new parameters (curve colours and styles matching Fig. 5), and the error $\delta_{\mathrm{sim}}-\delta_{\mathrm{HSW}}$ (lower panel), which is now visibly much smaller.
We interpret this as showing that our measure of effective radius is not interchangeable with that of others who fit HSW profiles. That is, our value $R_{\mathrm{v}}$ is more comparable to the meaning of “effective radius” as used in a $N$-body simulation or observations (e.g., Hamaus et al., 2014) than our $R_{\mathrm{eff}}$ is. This difference in notions of effective radius may be a resolution effect, due to our use of the continuous fluid approximation. The large physical size of the grid cells as compared to structures means that thinnest possible wall in the simulations is much thicker than a wall observed in data or one in an $N$-body simulation. This is supported by the fact that our best-fit $R_{\mathrm{v}}$ values are converging towards $R_{\mathrm{eff}}$ as we increase physical resolution (Table 1). Note that a further increase in resolution is unachievable in practice due to the limits of the fluid approximation. This exclusion of over-dense cells from voids may be another cause for this difference, as it causes the peak in density to occur further from $r=R_{\mathrm{eff}}$. For a direct comparison to results from Newtonian $N$-body simulations – for which the HSW profiles provide a very good fit – we would need to simulate collisionless particles alongside NR.
@@
4 $h^{-1}$ Mpc, using $R_{\mathrm{eff}}$0.668 1.02 $R_{\mathrm{eff}}$ 5.66 8.05
4 $h^{-1}$ Mpc, fitting $R_{\mathrm{v}}$0.690 1.06 $1.25\,R_{\mathrm{eff}}$ 3.24 8.10
12 $h^{-1}$ Mpc, using $R_{\mathrm{eff}}$0.405 1.05 $R_{\mathrm{eff}}$ 3.23 6.60
12 $h^{-1}$ Mpc, fitting $R_{\mathrm{v}}$0.426 1.10 $1.32\,R_{\mathrm{eff}}$ 2.03 9.38
In light of this, it may be more meaningful to rescale the void size functions of Section 4.1 such that the “effective radius” giving the “size” of the void is $R_{\mathrm{v}}=1.25R_{\mathrm{eff}}$ or $1.32R_{\mathrm{eff}}$ (Table 1) rather than the effective radius $R_{\mathrm{eff}}$ itself as we define it in (4), or in some other way compensate for the decrease in volume that results from low physical resolution.
@@ 4.4.4 General-relativistic radial profiles
Figure 7: The average fluid-intrinsic expansion $\Theta$ on a spherical shell at distance $r$ from the centre of volume, with each void scaled by its effective radius $R_{\mathrm{eff}}$. The shaded areas show the middle  68.5% and  95% of values for expansion on that shell. A horizontal line shows the whole-box average expansion $\Theta=3\bar{H}$. The upper panel shows the 4 $h^{-1}$ Mpc resolution simulation, the lower, 12 $h^{-1}$ Mpc.
Figure 8: Upper: The average fluid-intrinsic curvature scalar $\tensor\tensor[]{\mathcal{R}}{}$ on a spherical shell at distance $r$ from the centre of volume, with each void scaled by its effective radius $R_{\mathrm{eff}}$. The shaded areas show the middle  68.5% and  95% of values for the Ricci scalar on that shell. The upper panel shows the 4 $h^{-1}$ Mpc resolution simulation, the lower, 12 $h^{-1}$ Mpc.
Figure 9: $\Omega_{\mathrm{K}}$ averaged over shells of radius $r$ centred on 2000 stacked void centres of volume, each void scaled by its effective radius $R_{\mathrm{eff}}$. The solid curve shows the 4 $h^{-1}$ Mpc resolution simulation, the dashed curve, 12 $h^{-1}$ Mpc.
We also produce stacked radial profiles for three general-relativistic scalars, which are important in the Buchert averaging scheme (BMR). To obtain dimensionless profiles (analogous to obtaining a profile for $\delta$) we normalise the expansion and curvature scalar by the effective Hubble rate. The effective Hubble function $H_{\mathcal{D}}$ (over a spatial domain $\mathcal{D}$) is defined as
$H_{\mathcal{D}}=\frac{\dot{a}_{\mathcal{D}}}{a_{\mathcal{D}}},$ (10)
where the effective scale factor $a_{\mathcal{D}}$ is defined in terms of the evolution of the fluid volume $\mathcal{V}_{\mathcal{D}}(t)$ compared to its initial volume $\mathcal{V}_{\mathcal{D}_{i}}$
@@
$\dot{a}_{\mathcal{D}}=\mathrm{d}a_{\mathcal{D}}/\mathrm{d}t$, and the volume is defined in (3). The average scale factor over the entire simulation volume is denoted $\bar{a}$, and $\bar{H}=\dot{\bar{a}}/\bar{a}$. These coincide with expressions used in (1).
The first scalar we study is the expansion scalar in the rest frame of the fluid. $\Theta\equiv\nabla_{\mu}u^{\mu}$, where $\nabla_{\mu}$ is the covariant derivative associated with $g_{\mu\nu}$ and $u^{\mu}$ is the 4-velocity of the fluid. We calculate $\Theta$ (and $\bar{H}$ and the Ricci scalar) using existing routines in mescaline (Macpherson et al., 2019).
@@
This explicitly shows that voids expand at least $\sim$10–30faster than the large-scale average, as expected. The average expansion rate is lower on shells that intersect the walls, to a minimum shell average of 91% and 97% (4 $h^{-1}$ Mpc resolution and 12 $h^{-1}$ Mpc, respectively) of the large-scale average. In some cases individual cells have negative $\Theta$, implying collapse. Our expansion profiles here can be considered as lower limits on the expansion in the centre and walls of voids, due to the coarse smoothing scales in our simulations.
We also study the fluid-intrinsic curvature scalar $\tensor\tensor[]{\mathcal{R}}{}$ (defined in equation (4.15) of BMR). This can be roughly considered the 3-Ricci scalar in the rest frame of the fluid. Normalising this by $6\bar{H}^{2}$ gives a dimensionless quantity. Fig. 8 shows the stacked spherical profiles of $\tensor\tensor[]{\mathcal{R}}{}$ (solid curve) in the simulations at 4 $h^{-1}$ Mpc (upper panel) and 12 $h^{-1}$ Mpc (lower panel) resolution. In both panels, the shaded regions show the middle  68.5% and  95% of $\tensor\tensor[]{\mathcal{R}}{}$ values on each spherical shell.
We see that voids in our simulations have negative spatial curvature, while in the walls the spatial curvature is positive. As the distance from the void increases beyond the wall, the average spatial curvature tends towards the whole-volume average spatial curvature (flat; see Macpherson et al., 2019).
@@
though we note that unlike Buchert, we do not perform any spatial averaging of the curvature; only stacking of void profiles. For our radial profiles, the domain is each spherical shell, so $\Omega_{\mathrm{K}}$ is calculated using the values of $\Theta$ and $\tensor\tensor[]{\mathcal{R}}{}$ from the stacked spherical profiles above. We define $\bar{\Omega}_{\mathrm{K}}$ in (1) using (12) with the entire box as the domain.
Fig. 9 shows the average value of $\Omega_{\mathrm{K}}$. The magnitude of $\Omega_{\mathrm{K}}$ reaches 0.8 in the void centre for the smallest physical resolution we study here. We expect this to be even larger for a situation with resolution below 4 $h^{-1}$ Mpc (i.e., an $N$-body simulation or galaxy data). While the negative curvature (Fig. 9 and Fig. 8) and higher expansion (Fig. 7) in void centres are expected in general relativity, this work is the first time it has been shown explicitly using realistic simulations in NR (though see Macpherson et al., 2018).

### v2

@@ 4.2 Void fraction
Voids comprise the majority of the volume of the late-time Universe. Constraining the precise fraction of voids is important for the timescape model of cosmic backreaction (Wiltshire, 2007a, b, 2009; Duley et al., 2013), since it is the free parameter which replaces $\Omega_{\rm M}$ in standard FLRW models.
@@ 4.3 Finding void centres
We consider two choices of $f(\rho)$ which ensure that the centre is weighted towards areas of low density, namely $f(\rho)=1/\rho$ and $f(\rho)=\rho_{\mathrm{ave}}-\rho$. When matter is distributed unevenly around a void, its density-weighted centre will differ from its centre of volume. Under these definitions, the decision to include or exclude a higher-density cell has less impact on the centre that it would have on the centre of volume. Thus, it may be more robust to the choice of exclusion threshold or slight changes in the density near it, which may cause the set of excluded cells to change.
In the 4 $h^{-1}$ Mpc resolution simulation, the mean distance between a void’s density-weighted centre $\mathbf{X}^{\mathrm{cm}}_{v}$ and its centre of volume $\mathbf{X}^{\mathrm{bc}}_{v}$ is approximately 0.78 $h^{-1}$ Mpc, or  0.20 grid cells, for $f(\rho)=1/\rho$, and 0.88 $h^{-1}$ Mpc (0.22 grid cells) when $f(\rho)=\rho_{\mathrm{ave}}-\rho$. In both cases the difference has an upper bound of a little over one cell. The mean distance between the two definitions of $\mathbf{X}^{\mathrm{cm}}_{v}$ is 0.27 $h^{-1}$ Mpc. Since these differences are small with respect to our grid resolution, we consider all three to be appropriates measures of centre in our case. We use the inverse-density weighted centre $\mathbf{X}^{\mathrm{cm}}_{v}$, $f(\rho)=1/\rho$, hereafter for the sake of comparison with existing literature.
@@ 4.4 Radial profiles
Figure 3: Upper: The stacked density contrast (solid curve) on a spherical shell at distance $r$ from the stacked void centre in the 4 $h^{-1}$ Mpc (left) and 12 $h^{-1}$ Mpc (right) simulations, with each void individually normalised by its effective radius $R_{\mathrm{eff}}$. The shaded areas show the middle  68.5% and  95% of $\delta$ values for cells on that shell. Lower: The fraction of cells inside the void as a function of radius (solid) compared to perfect spheres (dotted) and a set of grid-approximated spheres with radius 3–5 grid cells (dashed).
@@ 4.4.1 Constructing stacked radial profiles
Because the steps are in terms of a fraction of each void’s radius, individual profiles can easily be stacked and averaged to produce a radial profile for a sample of voids. The process of stacking a large number of voids produces a spherically symmetric object, despite individual voids not being necessarily spherical. Stacking voids in this manner has other uses beyond extracting radial profiles. It can also be used to amplify the signal of the integrated Sachs–Wolfe effect (Granett et al., 2008), as temperature fluctuations in the CMB are larger than the temperature change due to these effects in any single structure; and for the Alcock–Paczyński (1979) test (Sutter et al., 2014a; Hamaus et al., 2020).
@@ 4.4.2 Stacked radial density profiles
We stack all of the voids in the simulation to produce averaged radial profiles for the density contrast in the voids. The top panels of Fig. 3 show these profiles, normalised by each void’s effective radius, for the 4 $h^{-1}$ Mpc and 12 $h^{-1}$ Mpc resolution simulations, respectively. The shaded areas of these figures show the density contrast of the middle  68.5% and  95% of cells at each rescaled radial distance. These shaded bands show that walls begin appearing in some directions even when $r<R_{\mathrm{eff}}$, where the average density contrast is still negative. Likewise, the overcompensation continues much further away from the centre than $R_{\mathrm{eff}}$, continuing to be visible past $r=2R_{\mathrm{eff}}$.
This is a natural consequence of the fact that the voids are not all close to spherical. The lower panels in these figures show this explicitly, showing what proportion of points at a given distance are inside of the void or outside of it, either in the walls or in a different void (solid curve). For comparison, these panels also show the case of spheres at infinite resolution (dotted line; a step function) and a set of spheres approximated on the grid with radius 3–5 grid cells (dashed curve) – comparable to the effective radius of typical voids in these simulations. The latter is not a step function, due to the spheres being represented on the finite-resolution grid. However, the transition between cells inside and outside of the void occurs over a shorter range of radial distances for spheres on the grid as compared to the simulated voids, confirming that the simulated voids are generally not spherical.
@@ 4.4.3 Fitting HSW profiles
Figure 4: The stacked radial density contrast profiles from Fig. 3 and the corresponding fitted HSW profiles. Namely, the 4 $h^{-1}$ Mpc simulation density contrast (solid blue), its fitted HSW profile (dashed black), the 12 $h^{-1}$ Mpc simulation density contrast (dash-dotted orange), and its fitted HSW profile (dotted black). These HSW profiles use (8), fitting only the typical four parameters.
Figure 5: Upper: The stacked radial density contrast profiles from Fig. 3 and the corresponding fitted HSW profiles. Namely, the 4 $h^{-1}$ Mpc simulation density contrast (solid blue), its fitted HSW profile (dashed black), the 12 $h^{-1}$ Mpc simulation density contrast (dash-dotted orange), and its fitted HSW profile (dotted black). These HSW profiles use (9), with the fifth fitted parameter $R_{\mathrm{v}}$ in place of $R_{\mathrm{eff}}$ to determine the profile. The $x$-axis is normalised by the original $R_{\mathrm{eff}}$. Lower: Difference between HSW density contrast and stacked radial density contrast for the 4 $h^{-1}$ Mpc resolution simulation (solid) and the 12 $h^{-1}$ Mpc resolution simulation (dash-dotted).
@@
where $R_{\mathrm{v}}$ is a parameter that determines the size of the void in place of $R_{\mathrm{eff}}$. The rows of Table 1 with “fitting $R_{\mathrm{v}}$” list the adjusted parameter sets. The fitted $R_{\mathrm{v}}$ is 25% and 34% larger than the effective radius for 4 $h^{-1}$ Mpc and 12 $h^{-1}$ Mpc resolution, respectively. Fig. 5 shows the HSW profiles with these new parameters (curve colours and styles matching Fig. 4), and the error $\delta_{\mathrm{sim}}-\delta_{\mathrm{HSW}}$ (lower panel), which is now visibly much smaller.
@@
4 $h^{-1}$ Mpc, using $R_{\mathrm{eff}}$0.673 1.02 $R_{\mathrm{eff}}$ 5.79 8.02
4 $h^{-1}$ Mpc, fitting $R_{\mathrm{v}}$0.693 1.06 $1.25\,R_{\mathrm{eff}}$ 3.31 7.85
12 $h^{-1}$ Mpc, using $R_{\mathrm{eff}}$0.400 1.02 $R_{\mathrm{eff}}$ 3.39 6.60
12 $h^{-1}$ Mpc, fitting $R_{\mathrm{v}}$0.424 1.08 $1.34\,R_{\mathrm{eff}}$ 2.01 9.17
We interpret this as showing that our measure of effective radius is not interchangeable with that of others who fit HSW profiles. That is, our value $R_{\mathrm{v}}$ is more comparable to the meaning of “effective radius” as used in a $N$-body simulation or observations (e.g., Hamaus et al., 2014) than our $R_{\mathrm{eff}}$ is. This difference in notions of effective radius may be a resolution effect, due to our use of the continuous fluid approximation. The large physical size of the grid cells as compared to structures means that thinnest possible wall in the simulations is much thicker than a wall observed in data or one in an $N$-body simulation. This is supported by the fact that our best-fit $R_{\mathrm{v}}$ values are converging towards $R_{\mathrm{eff}}$ as we increase physical resolution (Table 1). Note that a further increase in resolution is unachievable in practice due to the limits of the fluid approximation.
This exclusion of over-dense cells from voids may be another cause for this difference, as it causes the peak in density to occur further from $r=R_{\mathrm{eff}}$. For a direct comparison to results from Newtonian $N$-body simulations – for which the HSW profiles provide a very good fit – we would need to simulate collisionless particles alongside NR.
In light of this, it may be more meaningful to rescale the void size functions of Section 4.1 such that the “effective radius” giving the “size” of the void is $R_{\mathrm{v}}=1.25R_{\mathrm{eff}}$ or $1.34R_{\mathrm{eff}}$ (Table 1) rather than the effective radius $R_{\mathrm{eff}}$ itself as we define it in (4), or in some other way compensate for the decrease in volume that results from low physical resolution.
@@ 4.4.4 General-relativistic radial profiles
Figure 6: The average fluid-intrinsic expansion $\Theta$ on a spherical shell at distance $r$ from the stacked void centres, with each void scaled by its effective radius $R_{\mathrm{eff}}$. The shaded areas show the middle  68.5% and  95% of values for expansion on that shell. A horizontal line shows the whole-box average expansion $\Theta=3\bar{H}$. The upper panel shows the 4 $h^{-1}$ Mpc resolution simulation, the lower, 12 $h^{-1}$ Mpc.
Figure 7: Upper: The average fluid-intrinsic curvature scalar $\tensor\tensor[]{\mathcal{R}}{}$ on a spherical shell at distance $r$ from the stacked void centres, with each void scaled by its effective radius $R_{\mathrm{eff}}$. The shaded areas show the middle  68.5% and  95% of values for the Ricci scalar on that shell. The upper panel shows the 4 $h^{-1}$ Mpc resolution simulation, the lower, 12 $h^{-1}$ Mpc.
Figure 8: Median $\Omega_{\mathrm{K}}$ over all voids for each distance $r/R_{\mathrm{eff}}$. For each void, the domain of $\Omega_{\mathrm{K}}$ is a shell of radius $r$ centred on the void’s centre, with the void scaled by its effective radius $R_{\mathrm{eff}}$. The solid curve shows the median for voids in the 4 $h^{-1}$ Mpc resolution simulation, the dashed curve, 12 $h^{-1}$ Mpc.
We also produce stacked radial profiles for three general-relativistic scalars, which are important in the Buchert averaging scheme (BMR). To obtain dimensionless profiles (analogous to obtaining a profile for $\delta$) we normalise the expansion and curvature scalar by the effective Hubble rate. The effective Hubble function $H_{\mathcal{D}}$ (over a spatial domain $\mathcal{D}$) is defined in terms of the average expansion as
$H_{\mathcal{D}}:=\frac{\dot{a}_{\mathcal{D}}}{a_{\mathcal{D}}}=\frac{1}{3}\langle\Theta\rangle_{\mathcal{D}}\,.$ (10)
Thus the effective scale factor $a_{\mathcal{D}}$ is defined in terms of the evolution of the fluid volume $\mathcal{V}_{\mathcal{D}}(t)$ compared to its initial volume $\mathcal{V}_{\mathcal{D}_{i}}$
@@
$\dot{a}_{\mathcal{D}}=\mathrm{d}a_{\mathcal{D}}/\mathrm{d}t$, and the volume is defined in (3). The average scale factor over the entire simulation volume is denoted $\bar{a}$, and $\bar{H}=\bar{\Theta}/3$.
The first scalar we study is the expansion scalar in the rest frame of the fluid, $\Theta\equiv\nabla_{\mu}u^{\mu}$, where $\nabla_{\mu}$ is the covariant derivative associated with $g_{\mu\nu}$ and $u^{\mu}$ is the 4-velocity of the fluid. We calculate $\Theta$ (and $\bar{H}$ and the Ricci scalar) using existing routines in mescaline (Macpherson et al., 2019).
@@
This explicitly shows that voids expand at least $\sim$10–30faster than the large-scale average. This is in agreement with predictions from linear perturbation theory (e.g., Lahav et al., 1991) and previous NR studies (e.g., Fig. 10 of Macpherson et al., 2019). The average expansion rate is lower on shells that intersect the walls, to a minimum shell average of 91% and 97% (4 $h^{-1}$ Mpc resolution and 12 $h^{-1}$ Mpc, respectively) of the large-scale average. In some cases individual cells have negative $\Theta$, implying collapse. Our expansion profiles here can be considered as lower limits on the expansion in the centre and walls of voids, due to the coarse smoothing scales in our simulations.
We also study the fluid-intrinsic curvature scalar $\tensor\tensor[]{\mathcal{R}}{}$ (defined in equation (4.15) of BMR). This can be roughly considered the 3-Ricci scalar in the rest frame of the fluid. Normalising this by $6\bar{H}^{2}$ gives a dimensionless quantity. Fig. 7 shows the stacked spherical profiles of $\tensor\tensor[]{\mathcal{R}}{}$ (solid curve) in the simulations at 4 $h^{-1}$ Mpc (upper panel) and 12 $h^{-1}$ Mpc (lower panel) resolution. In both panels, the shaded regions show the middle  68.5% and  95% of $\tensor\tensor[]{\mathcal{R}}{}$ values on each spherical shell. We see that voids in our simulations have negative spatial curvature, while in the walls the spatial curvature is positive. As the distance from the void increases beyond the wall, the average spatial curvature tends towards the whole-volume average spatial curvature (flat; see Macpherson et al., 2019).
@@
Buchert introduces a general formalism for spatial averages on arbitrary domains, $\mathcal{D}$, which leads to markedly different results depending on the domain of averaging. To study the effect of small scale inhomogeneities, the domain $\mathcal{D}$ is that of our fluid elements as limited by the resolution scale of the simulation. The procedure of stacking voids over the entire box is then an operational realisation of how the volume average void fraction is defined in the timescape interpretation of the Buchert scheme (Wiltshire, 2007a, 2009).
For our radial profiles, the averaging domain is an individual spherical shell. Thus, we find a $\Omega_{\mathrm{K}}$ profile for a single void by finding the average values $\Theta$ and $\tensor\tensor[]{\mathcal{R}}{}$ on individual shells of radius $r$ around that void. The profile for each void is then rescaled by the void’s effective radius $R_{\mathrm{eff}}$ and the median $\Omega_{\mathrm{K}}$ of the set of all void profiles is found at each distance $r/R_{\mathrm{eff}}$ to produce the stacked profile, shown in Fig. 8.
The magnitude of $\Omega_{\mathrm{K}}$ reaches 0.8 in the void centre for the smallest physical resolution we study here. We expect this to be even larger for a situation with resolution below 4 $h^{-1}$ Mpc (i.e., an $N$-body simulation or galaxy data). While the negative curvature (Fig. 8 and Fig. 7) and higher expansion (Fig. 6) in void centres are expected in GR, this work is the first time it has been shown explicitly using realistic simulations in NR (though see Macpherson et al., 2018).